# Retrosynthesis Planning via Worst-path Policy Optimisation in Tree-structured MDPs

### v1

@@ 5.2 Ablation Studies
We proposed a value function to estimate the worst-path return. The worst-path return indicates the estimated depth of the synthesis tree: $\text{depth}(s)=\frac{\log V_{\phi}(s)}{\log\gamma}$. In Figure 4(b), we investigate the distribution of the estimated depth, corresponding to the real depth by direct generation. We found that the value function can reflect the difficulties to synthesise the molecules, while it shows better capability on the molecules that require a less deep synthetic route.
Advantage Coefficient.
@@
The hyperparameter $\beta$ controls the imitation strength on the past successful experience. A large $\beta$ means a focused imitation on the action with a high advantage. Figure 4(c) shows the learning curve on the Retro*-190 test set with $\beta\in\{0,1,2,5,10,20\}$. The figure shows that the DG performance increases from $16\%$ to around $80\%$ by uniformly imitating successful past experiences. With the advantage weighting, the performance soars to a success rate of more than $90\%$ when $\beta$ increases to $2$, and more than $95\%$ with $\beta\geq 10$.
6 Discussion and Conclusions
@@
We introduced InterRetro, a novel approach that frames retrosynthesis planning as worst-path optimisation in tree-structured MDPs. Our weighted self-imitation algorithm enables single-step models to generate high-quality synthetic routes without search at inference time, achieving SOTA performance in success rate, route length, and sample efficiency. Whilst this worst-path policy optimisation has proven effective, future investigations could explore alternative worst-path criteria, such as minimising the maximum depth among all successful paths, incorporating the cost of the reaction and the building blocks, or designing a heuristic search to optimise the worst-path objective.
@@
Whilst our formulation offers significant advantages, it assumes that all proposed reactions are feasible in real-world settings — an assumption that does not always hold. Although we mitigate this by imitating only actions supported by pretrained models (similar to [1, 10]), computational approaches cannot guarantee practical feasibility without experimental validation. Real chemical reactions depend on specific conditions that are rarely captured in retrosynthesis datasets. Future work should therefore focus on creating related datasets and integrating reaction feasibility checks into InterRetro. Moreover, this “weakest link” principle is broadly applicable beyond chemistry to sequential decision problems where overall success depends on the weakest component, such as robust project planning where delays in any critical task affect the entire timeline, and multi-agent systems where team performance is limited by the least capable agent.
Appendix A Illustrative Example: Worst-Path vs. Best-Path Optimisation
To further clarify the motivation behind our worst-path objective, we present a simplified pedagogical example that illustrates the critical “weakest link” principle in retrosynthesis planning. Using the same tree-structured MDP framework (Section 3.1), we demonstrate why optimising for the worst-performing path is essential for practical synthetic route design.
Consider a scenario with discount factor $\gamma=0.9$ and a target molecule $M_{0}$ that can be addressed by two distinct initial reactions, $a_{1}$ and $a_{2}$. Each reaction produces reactants that must all be successfully processed for the overall route to be viable. We define two action-value functions for a state $s$ and action $a$:
$J_{WP}(s,a)$ $=\min_{p\in P(\mathcal{T}(s,a))}V(p),\quad$ (Worst-Path Action-Value) (17)
$J_{BP}(s,a)$ $=\max_{p\in P(\mathcal{T}(s,a))}V(p),\quad$ (Best-Path Action-Value) (18)
where $P(\mathcal{T}(s,a))$ is the set of all root-to-leaf paths in the tree generated by taking action $a$ from state $s$, and $V(p)$ is the discounted return of a specific path $p$. The Worst-Path (WP) objective evaluates actions by their worst-case outcome, whilst the Best-Path (BP) objective considers only the most favourable outcome.
$M_{0}$ $R_{1A}$ $R_{1B}$ $BB_{1}$ $E$ $a_{1}$ Tree $T_{1}$: $J_{WP}(M_{0},a_{1})=0$ $J_{BP}(M_{0},a_{1})=0.81$ $M_{0}$ $R_{2A}$ $R_{2B}$ $R^{\prime}_{2A}$ $R^{\prime}_{2B}$ $B_{2} BB$$B_{3}$ $a_{2}$ Tree $T_{2}$: $J_{WP}(M_{0},a_{2})=0.729$ $J_{BP}(M_{0},a_{2})=0.729$
Figure 5: An illustrative Tree MDP comparing two potential reaction pathways from target molecule $M_{0}$ (blue circles). Each reaction (action $a_{1}$ or $a_{2}$) decomposes $M_{0}$ into multiple reactants that must all be synthesised. Yellow circles labelled $BB_{i}$ represent purchasable building blocks that yield reward 1, while the red circle labelled $E$ represents a dead-end with reward 0. Tree $T_{1}$ (left) has one successful 2-step path and one failed path. Tree $T_{2}$ (right) has two successful 3-step paths. Path values $V(p)$ are discounted by $\gamma=0.9$, giving $0.9^{2}=0.81$ for 2-step paths and $0.9^{3}=0.729$ for 3-step paths.
Let $V(p)$ be the discounted return of a complete path $p$ starting from $M_{0}$ and ending at a leaf node, calculated as $\gamma^{L}r(\text{leaf})$, where $L$ is the total path length from $M_{0}$, and $r(\text{leaf})$ is 1 for a $BB$ and 0 for a dead-end $E$.
Evaluating Action $a_{1}$ (leading to Tree $T_{1}$):
  • Path $p_{1A}$ (via $R_{1A}$ to $BB_{1}$): Length $L=2$. $V(p_{1A})=\gamma^{2}\cdot 1=(0.9)^{2}=0.81$.
  • Path $p_{1B}$ (via $R_{1B}$ to $E$): Length $L=2$. $V(p_{1B})=\gamma^{2}\cdot 0=0$.
The worst-path action-value for $a_{1}$ is:
$J_{WP}(M_{0},a_{1})=\min\{V(p_{1A}),V(p_{1B})\}=\min\{0.81,0\}=0.$
Evaluating Action $a_{2}$ (leading to Tree $T_{2}$):
  • Path $p_{2A}$ (via $R_{2A}$, $R^{\prime}_{2A}$ to $BB_{2}$): Length $L=3$. $V(p_{2A})=\gamma^{3}\cdot 1=(0.9)^{3}=0.729$.
  • Path $p_{2B}$ (via $R_{2B}$, $R^{\prime}_{2B}$ to $BB_{3}$): Length $L=3$. $V(p_{2B})=\gamma^{3}\cdot 1=(0.9)^{3}=0.729$.
The worst-path action-value for $a_{2}$ is:
$J_{WP}(M_{0},a_{2})=\min\{V(p_{2A}),V(p_{2B})\}=\min\{0.729,0.729\}=0.729.$
Therefore, a policy $\pi_{WP}$ optimising the worst-path objective (Eq. 17) would choose action $a_{2}$ (value $0.729$) over action $a_{1}$ (value 0).
Now, considering the Best-Path Seeking objective (Eq. 18):
  • For action $a_{1}$: $J_{BP}(M_{0},a_{1})=\max\{V(p_{1A}),V(p_{1B})\}=\max\{0.81,0\}=0.81$.
  • For action $a_{2}$: $J_{BP}(M_{0},a_{2})=\max\{V(p_{2A}),V(p_{2B})\}=\max\{0.729,0.729\}=0.729$.
A policy $\pi_{BP}$ optimising this best-path objective would choose action $a_{1}$ (value $0.81$) over action $a_{2}$ (value $0.729$).
This example highlights a key difference: the Best-Path policy $\pi_{BP}$ is drawn to the higher potential reward of path $p_{1A}$ in $T_{1}$, despite the significant risk of complete failure associated with path $p_{1B}$. In contrast, our worst-path policy $\pi_{WP}$ strictly prefers the entirely viable option $a_{2}$ (albeit with individually longer paths), ensuring that all branches of the synthesis lead to success. This inherent risk aversion is crucial for practical retrosynthesis planning, where a single unresolvable intermediate invalidates the entire proposed route.
Appendix B Case Studies
@@
In Figure 6, we randomly select three targets from the Retro*-190 benchmark and illustrate their complete synthesis trees as directly generated by InterRetro in a single forward pass. (i) For the sulfonylated chiral cyclohexanol derivative, InterRetro proposes a four-step route involving: a sulfonylation of a chiral alcohol to introduce the sulfonate ester, an electrophilic bromination of the diazine core, a C(sp3)–C(sp2) cross-coupling to build the fused triazolopyrimidine system, and an N-arylation to complete the final structure. (ii) For the alkyne-substituted diazine, InterRetro constructs the target via two steps: first an acylation of the brominated azine to give the N-formyl intermediate, followed by an SN2 N-propargylation. (iii) For the complex triazolopyridine-fused drug-like scaffold, InterRetro suggests a three-step route: a benzylic etherification that assembles the protected piperidine core, a C(sp³)–C(sp²) cross-coupling to append the triazolopyridine fragment, and an N-alkylation that installs the silyl-protected side chain. Across all three cases, InterRetro applies feasible chemical transitions and terminates in commercially available building blocks. A full list of the synthetic routes on Retro*-190 is attached in the supplementary materials.
Appendix C Missing Proofs and Further Theory
C.1 Proof of Proposition C.1
@@ C.2 Proof of Proposition 2
C.3 Bellman Optimality Operator and Contraction Mapping
@@ C.4 Proof of Proposition 4
where $Z(s)=\sum_{a^{\prime}}\pi^{i}(a^{\prime}\mid s)\cdot\exp(\beta A^{\pi^{i}}(s,a^{\prime}))$ is a normalizing constant. This update ensures that $\pi^{i+1}$ remains in the same support set as $\pi^{i}$.
@@
Appendix D Graph2Edits as Single-step Model
Graph2Edits [38] considers one-step retrosynthesis as a graph-editing game. Starting from the product molecule, the model autoregressively predicts a short sequence of primitive edits — Delete Bond, Change Bond, Change Atom, Attach Leaving Group, and finally a Terminate token. At each step the current intermediate graph is embedded by a directed message-passing neural network (D-MPNN), whose atom- and bond-level embeddings are fed to three linear heads that score every possible bond edit, atom edit and the termination symbol. The highest-scoring edit is applied to yield the next intermediate, and the process repeats until Terminate is chosen, at which point the intermediates have been fully converted to the predicted reactants. This edit vocabulary, mined from training data, covers $99.9\%$ of USPTO-50k reactions and encodes stereochemistry as well as functional leaving groups.
Trained with teacher forcing, the system already achieves a $55\%$ top-1 exact-match accuracy on USPTO-50k and maintains high validity and diversity even for long or stereochemically rich reactions; these strengths make it an ideal single-step model inside our multi-step InterRetro framework.
Appendix E Broader Impacts
By generating complete retrosynthetic routes in a single forward pass, InterRetro can shorten early-stage discovery cycles for drug discovery and materials design. The method’s tendency to propose shorter, higher-yielding sequences also points toward reduced solvent usage, energy consumption, and hazardous waste, aligning with global goals for greener and more sustainable chemistry. However, a search-free planner lowers technical barriers for malicious actors who might design routes to controlled substances or chemical weapons, and the model’s assumption of universal feasibility could mislead non-experts into attempting unsafe or impractical experiments. Therefore, we advise all users to carefully weigh the positive and negative societal impacts before employing InterRetro, and to pair any deployment with rigorous safety and feasibility checks.

### v2

@@ 5.2 Ablation Studies
Value Estimation. We proposed a value function to estimate the worst-path return. The worst-path return indicates the estimated depth of the synthetic tree: $\text{depth}(s)=\frac{\log V_{\phi}(s)}{\log\gamma}$. In Figure 4(b), we investigate the distribution of the estimated depth, corresponding to the real depth by direct generation. We found that the value function can reflect the difficulties to synthesise the molecules, while it shows better capability on the molecules that require a less deep synthetic route.
@@
Advantage Coefficient. The hyperparameter $\beta$ controls the imitation strength on the past successful experience. A large $\beta$ means a focused imitation on the action with a high advantage. Figure 4(c) shows the learning curve on the Retro*-190 test set with $\beta\in\{0,1,2,5,10,20\}$. The figure shows that the DG performance increases from $16\%$ to around $80\%$ by uniformly imitating successful past experiences. With the advantage weighting, the performance soars to a success rate of more than $90\%$ when $\beta$ increases to $2$, and more than $95\%$ with $\beta\geq 10$.
6 Conclusions and Discussion
@@
We introduced InterRetro, a novel approach that frames retrosynthesis planning as worst-path optimisation in tree-structured MDPs. Our weighted self-imitation algorithm enables single-step models to generate high-quality synthetic routes without search at inference time, achieving SOTA performance in success rate, route length, and sample efficiency.
@@
Further investigation can proceed in three main directions. First, the worst-path objective could be extended to incorporate additional real-world constraints, such as reaction conditions and the cost of building blocks. Second, and perhaps most importantly, our method and other contemporary work assume that all proposed reactions are feasible in real-world settings — an assumption that does not always hold. Although we mitigate this by imitating only actions supported by pre-trained models, computational approaches cannot guarantee practical feasibility without experimental validation. Future work can therefore focus on integrating reaction feasibility checks into InterRetro. Furthermore, the “weakest-link” principle underlying our method is broadly applicable beyond chemistry, extending to sequential decision problems where overall success depends on the least reliable component — for example, in robust project planning where delays in any critical task impact the entire timeline, or in multi-agent systems where team performance is constrained by the least capable agent.
Broader Impacts. The retrosynthesis community has recently open-sourced many high-quality models capable of suggesting synthetic routes for a vast number of molecules — including both beneficial drugs and potentially harmful compounds. To promote safe and ethical research, we propose that access to such models or their source code be governed by a Responsible Use License, under which researchers acknowledge the responsible and lawful application of the technology.
Acknowledgments
We acknowledge support from a UKRI Turing AI Acceleration Fellowship (EPSRC EP/V024868/1).
Appendix A Case Studies
@@
In Figure 5, we randomly select three targets from the Retro*‑190 benchmark and illustrate their complete synthetic trees as directly generated by InterRetro in a single forward pass. (i) For the sulfonylated chiral cyclohexanol derivative, InterRetro proposes a four‑step route involving: a sulfonylation of a chiral alcohol to introduce the sulfonate ester, an electrophilic bromination of the diazine core, a C(sp3)–C(sp2) cross‑coupling to build the fused triazolopyrimidine system, and an N‑arylation to complete the final structure. (ii) For the alkyne‑substituted diazine, InterRetro constructs the target via two steps: first an acylation of the brominated azine to give the N‑formyl intermediate, followed by an SN2 N‑propargylation. (iii) For the complex triazolopyridine‑fused drug‑like scaffold, InterRetro suggests a three‑step route: a benzylic etherification that assembles the protected piperidine core, a C(sp³)–C(sp²) cross‑coupling to append the triazolopyridine fragment, and an N‑alkylation that installs the silyl‑protected side chain. Across all three cases, InterRetro applies feasible chemical transitions and terminates in commercially available building blocks. A full list of the synthetic routes on Retro*‑190 is attached in the supplementary materials.
Appendix B Proofs and Further Theory
B.1 Proof of Proposition 1
@@ B.2 Proof of Proposition 2
B.3 Proof of Proposition 3
@@ B.4 Proof of Proposition 4
where $Z(s)=\sum_{a^{\prime}}\pi^{i}(a^{\prime}\mid s)\cdot\exp(\beta A^{\pi^{i}}(s,a^{\prime}))$ is a normalising constant. This update ensures that $\pi^{i+1}$ remains in the same support set as $\pi^{i}$.
@@
Appendix C Graph2Edits as Single-step Model
Graph2Edits [zhong2023graph2edits] considers one‑step retrosynthesis as a graph‑editing game. Starting from the product molecule, the model autoregressively predicts a short sequence of primitive edits — Delete Bond, Change Bond, Change Atom, Attach Leaving Group, and finally a Terminate token. At each step the current intermediate graph is embedded by a directed message‑passing neural network (D‑MPNN), whose atom‑ and bond‑level embeddings are fed to three linear heads that score every possible bond edit, atom edit and the termination symbol. The highest‑scoring edit is applied to yield the next intermediate, and the process repeats until Terminate is chosen, at which point the intermediates have been fully converted to the predicted reactants. This edit vocabulary, mined from training data, covers $99.9\%$ of USPTO‑50k reactions and encodes stereochemistry as well as functional leaving groups.
Trained with teacher forcing, the system already achieves a $55\%$ top‑1 exact‑match accuracy on USPTO‑50k and maintains high validity and diversity even for long or stereochemically rich reactions; these strengths make it an ideal single‑step model inside our multi‑step InterRetro framework.